# GVE-Louvain: Fast Louvain Algorithm for Community Detection in Shared Memory Setting

### v4

@@ 5.1. Experimental Setup
We use a server that has two $16$-core x86-based Intel Xeon Gold 6226R processors running at $2.90$ GHz. Each core has an L1 cache of $1$ MB, an L2 cache of $16$ MB, and a shared L3 cache of $22$ MB. The machine has $93.4$ GB of system memory and runs on CentOS Stream 8. We use GCC 8.5 and OpenMP 4.5. Table 2 shows the graphs we use in our experiments. All of them are obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019).
Table 2. List of $13$ graphs obtained SuiteSparse Matrix Collection (Kolodziej et al., 2019) (directed graphs are marked with *). Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges), $D_{avg}$ is the average degree, and $|\Gamma|$ is the number of communities obtained using Louvain algorithm.
@@
(a) Runtime in seconds (logarithmic scale) with Vite, Grappolo, NetworKit, and GVE-Louvain
(b) Speedup of GVE-Louvain with respect to Vite, Grappolo, NetworKit.
(c) Modularity of communities obtained with Vite, Grappolo, NetworKit, and GVE-Louvain.
Figure 5. Runtime in seconds (logarithmic scale), speedup, and modularity of communities obtained with Vite, Grappolo, NetworKit, and GVE-Louvain for each graph in the dataset.
@@ 5.2. Comparing Performance of GVE-Louvain
We now compare the performance of GVE-Louvain with Vite (Louvain) (Ghosh et al., 2018a), Grappolo (Louvain) (Halappanavar et al., 2017), and NetworKit Louvain (Staudt et al., 2016). For Vite, we convert the graph datasets to Vite’s binary graph format, run it on a single node with threshold cycling/scaling optimization, and measure the reported average total time. For Grappolo, we measure the run it on the same system, and measure the reported total time. For NetworKit, we use a Python script to invoke PLM (Parallel Louvain Method), and measure the total time reported with getTiming(). For each graph, we measure the runtime of each implementation five times, for averaging. We also record the modularity of communities obtained, as reported by each implementation.
Figure 5(a) shows the runtimes of Vite (Louvain), Grappolo (Louvain), NetworKit Louvain, and GVE-Louvain on each graph in the dataset. Figure 5(b) shows the speedup of GVE-Louvain with respect to each implementation mentioned above. GVE-Louvain is on average 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × faster than Vite, Grappolo, and NetworKit respectively. On the sk-2005 graph, GVE-Louvain finds communities in $6.8$ seconds, and thus achieve a processing rate of $560$ million edges/s. Figure 5(c) shows the modularity of communities obtained with each implementation. GVE-Louvain on average obtains $3.1\%$ higher modularity than Vite (especially on web graphs), and $0.6\%$ lower modularity than Grappolo and NetworKit (especially on social networks with poor clustering).
@@ 6. Conclusion
In conclusion, this report presents our parallel multicore implementation of the Louvain algorithm — a high quality community detection method, which, as far as we are aware, stands as the most efficient implementation of the algorithm on multicore CPUs. Here, we considered $9$ different optimizations, which significantly improve the performance of the local-moving and the aggregation phases of the algorithm. On a server with dual 16-core Intel Xeon Gold 6226R processors, comparison with competitive open source implementations (Vite and Grappolo) and packages (NetworKit) show that our optimized implementation of Louvain, which we term as GVE-Louvain, is on average 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × faster than Vite, Grappolo, and NetworKit respectively. In addition, GVE-Louvain on average obtains $3.1\%$ higher modularity than Vite, and $0.6\%$ lower modularity than Grappolo and NetworKit. On a web graph with $3.8$ billion edges, GVE-Louvain identifies communities in $6.8$ seconds, and thus achieves a processing rate of $560$ million edges/s. In addition, GVE-Louvain achieves a strong scaling factor of 1.6×1.6$\times$1.6 × for every doubling of threads. Looking ahead, future work could focus of designing dynamic algorithms for Louvain to accommodate dynamic graphs which evolve over time.
@@ Acknowledgements.
I would like to thank Prof. Kishore Kothapalli, Prof. Dip Sankar Banerjee, Balavarun Pedapudi, and Souvik Karfa for their support.
@@ A.1. Indirect Comparison with State-of-the-art Louvain Implementations
Chou and Ghosh (Chou and Ghosh, 2022) introduce Nido, a batched clustering method designed for GPUs. Their experimentation conducted on a system equipped with two $128$ core AMD EPYC 7742 CPUs running at $2.25$ GHz and eight NVIDIA A100 GPUs demonstrates a (geometric) mean speedup of 2.4×2.4$\times$2.4 × compared to Grappolo (Halappanavar et al., 2017), utilizing $128$ threads (as shown in Table 3 of their paper (Chou and Ghosh, 2022)). Additionally, they find that cuGraph (Hricik et al., 2020) offers a mean speedup of $\times7$ relative to Grappolo (also depicted in Table 3 of their paper). In contrast, our Louvain implementation achieves a mean speedup of $\times2$ compared to Grappolo. Consequently, it surpasses Chou and Ghosh by 9.2×9.2$\times$9.2 × and cuGraph by 2.9×2.9$\times$2.9 ×, again without employing a single GPU.
Gawande et al. (Gawande et al., 2022) introduce cuVite, an ongoing project focusing on distributed Louvain for heterogeneous systems. They evaluate cuVite, cuGraph, and Rundemanen on a single-GPU of NVIDIA DGX-2, i.e., NVIDIA Tesla V100, and with two $24$ core Intel Xeon Platinum 8168 CPUs running at $2.7$ GHz. Further, they evaluate Grappolo on a $224$ core system with eight $28$ core Intel Xeon Platinum 8276M CPUs running at $2.20$ GHz. They observe that, on average, cuGraph, Rundemanen, and cuVite exhibit speedups of $\times3$, 3.0×3.0$\times$3.0 ×, and 3.28×3.28$\times$3.28 × respectively compared to Grappolo. Consequently, our Louvain implementation achieves a speedup of 6.7×6.7$\times$6.7 × compared to both cuGraph and cuVite, and a speedup of 7.3×7.3$\times$7.3 × relative to Rundemanen, all achieved without the utilization of a GPU.

### v5

@@ 5.1. Experimental Setup
5.1.1. System used
We use a server that has two $16$-core x86-based Intel Xeon Gold 6226R processors running at $2.90$ GHz. Each core has an L1 cache of $1$ MB, an L2 cache of $16$ MB, and a shared L3 cache of $22$ MB. The machine has $93.4$ GB of system memory and runs on CentOS Stream 8. For our GPU experiments, we employ a system featuring an NVIDIA A100 GPU (108 SMs, 64 CUDA cores per SM, 80 GB global memory, 1935 GB/s bandwidth, 164 KB shared memory per SM) paired with an AMD EPYC-7742 processor (64 cores, 2.25 GHz). This server is equipped with 512 GB of DDR4 RAM and operates on Ubuntu 20.04.
5.1.2. Configuration
We employ 32-bit integers for vertex IDs and 32-bit floats for edge weights, but use 64-bit floats for both computations and hashtable values. Our implementation leverages $64$ threads to align with the number of cores on the system, unless otherwise specified. For compilation, we use GCC 8.5 and OpenMP 4.5 on the CPU system, while on the GPU system, we use GCC 9.4, OpenMP 5.0, and CUDA 11.4.
5.1.3. Dataset
The graphs utilized in our experiments are detailed in Table 2. These graphs are sourced from the SuiteSparse Matrix Collection (Kolodziej et al., 2019). The graphs have vertex counts ranging from $3.07$ to $214$ million and edge counts ranging from $25.4$ million to $3.80$ billion. We ensure that the edges are undirected and weighted, with a default weight of $1$.
Table 2. List of $13$ graphs obtained SuiteSparse Matrix Collection (Kolodziej et al., 2019) (directed graphs are marked with *). Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges), $D_{avg}$ is the average degree, and $|\Gamma|$ is the number of communities obtained using GVE-Louvain.
@@ 5.1.3. Dataset
(a) Runtime in seconds (logarithmic scale) with Vite (Louvain), Grappolo (Louvain), NetworKit Louvain, cuGraph Louvain, and GVE-Louvain
(b) Speedup of GVE-Louvain with respect to Vite (Louvain), Grappolo (Louvain), NetworKit Louvain, and cuGraph Louvain.
(c) Modularity of communities obtained with Vite (Louvain), Grappolo (Louvain), NetworKit Louvain, cuGraph Louvain, and GVE-Louvain.
Figure 5. Runtime in seconds (logarithmic scale), speedup, and modularity of communities obtained with Vite (Louvain), Grappolo (Louvain), NetworKit Louvain, cuGraph Louvain, and GVE-Louvain for each graph in the dataset.
@@ 5.2. Comparing Performance of GVE-Louvain
We now compare the performance of GVE-Louvain with Vite (Louvain) (Ghosh et al., 2018a), Grappolo (Louvain) (Halappanavar et al., 2017), NetworKit Louvain (Staudt et al., 2016), and cuGraph Louvain (Kang et al., 2023). For Vite, we convert the graph datasets to Vite’s binary graph format, run it on a single node with threshold cycling/scaling optimization, and measure the reported average total time. For Grappolo, we measure the run it on the same system, and measure the reported total time. For NetworKit Louvain, we use a Python script to invoke PLM (Parallel Louvain Method), and measure the total time reported with getTiming(). To test cuGraph’s Leiden algorithm, we write a Python script that configures the Rapids Memory Manager (RMM) to use a pool allocator for fast memory allocations. We then execute cugraph.leiden() on the loaded graph. For each graph, we measure the runtime and the modularity of the obtained communities (as reported by each implementation), performing five runs to calculate an average. When using cuGraph, we discard the runtime of the first run to ensure that subsequent measurements accurately reflect RMM’s pool usage without the overhead of initial CUDA memory allocation.
Figure 5(a) shows the runtimes of Vite (Louvain), Grappolo (Louvain), NetworKit Louvain, cuGraph Louvain, and GVE-Louvain on each graph in the dataset. cuGraph’s Leiden algorithm fails to run on the arabic-2005, uk-2005, webbase-2001, it-2004, and sk-2005 graphs because of out-of-memory issues. On the sk-2005 graph, GVE-Louvain finds communities in $6.8$ seconds, and thus achieve a processing rate of $560$ million edges/s. Figure 5(b) shows the speedup of GVE-Louvain with respect to each implementation mentioned above. GVE-Louvain is on average 50×50$\times$50 ×, $\times2$, 20×20$\times$20 ×, and 5.8×5.8$\times$5.8 × faster than Vite, Grappolo, NetworKit Louvain, and cuGraph Louvain respectively. Figure 5(c) shows the modularity of communities obtained with each implementation. GVE-Louvain on average obtains $3.1\%$ higher modularity than Vite (especially on web graphs), and $0.6\%$ lower modularity than Grappolo and NetworKit (especially on social networks with poor clustering), and $2.6\%$ higher modularity than cuGraph Louvain (primarily because cuGraph Leiden failed to run on graphs that are well-clusterable).
@@ 6. Conclusion
In conclusion, this report presents our parallel multicore implementation of the Louvain algorithm — a high quality community detection method, which, as far as we are aware, stands as the most efficient implementation of the algorithm on multicore CPUs. Here, we considered $9$ different optimizations, which significantly improve the performance of the local-moving and the aggregation phases of the algorithm. On a server with dual 16-core Intel Xeon Gold 6226R processors, comparison with competitive open source implementations (Vite and Grappolo) and packages (NetworKit) show that our optimized implementation of Louvain, which we term as GVE-Louvain, is on average 50×50$\times$50 ×, $\times2$, 20×20$\times$20 ×, and 5.8×5.8$\times$5.8 × faster than Vite, Grappolo, NetworKit Louvain, and cuGraph Louvain (run on an NVIDIA A100 GPU) respectively. In addition, GVE-Louvain on average obtains $3.2\%$ higher modularity than Vite, $0.6\%$ lower modularity than Grappolo and NetworKit Louvain, and $2.6\%$ higher modularity than cuGraph Louvain. On a web graph with $3.8$ billion edges, GVE-Louvain identifies communities in $6.8$ seconds, and thus achieves a processing rate of $560$ million edges/s. In addition, GVE-Louvain achieves a strong scaling factor of 1.6×1.6$\times$1.6 × for every doubling of threads. Future work could focus of designing fast community detection algorithms that enable interactive updation of community memberships of vertices in large dynamic graphs.
@@ Acknowledgements.
I would like to thank Prof. Kishore Kothapalli, Prof. Dip Sankar Banerjee, Balavarun Pedapudi, Souvik Karfa, Chuck Hastings, and Rick Ratzel for their support.
@@ A.1. Indirect Comparison with State-of-the-art Louvain Implementations
Chou and Ghosh (Chou and Ghosh, 2022) introduce Nido, a batched clustering method designed for GPUs. Their experimentation conducted on a system equipped with two $128$ core AMD EPYC 7742 CPUs running at $2.25$ GHz and eight NVIDIA A100 GPUs demonstrates a (geometric) mean speedup of 2.4×2.4$\times$2.4 × compared to Grappolo (Halappanavar et al., 2017), utilizing $128$ threads (as shown in Table 3 of their paper (Chou and Ghosh, 2022)). In contrast, our Louvain implementation achieves a mean speedup of $\times2$ compared to Grappolo. Consequently, it surpasses Chou and Ghosh by 9.2×9.2$\times$9.2 ×.
Gawande et al. (Gawande et al., 2022) introduce cuVite, an ongoing project focusing on distributed Louvain for heterogeneous systems. They evaluate cuVite, cuGraph, and Rundemanen on a single-GPU of NVIDIA DGX-2, i.e., NVIDIA Tesla V100, and with two $24$ core Intel Xeon Platinum 8168 CPUs running at $2.7$ GHz. Further, they evaluate Grappolo on a $224$ core system with eight $28$ core Intel Xeon Platinum 8276M CPUs running at $2.20$ GHz. They observe that, on average, Rundemanen and cuVite exhibit speedups of 3.0×3.0$\times$3.0 × and 3.28×3.28$\times$3.28 × respectively compared to Grappolo. Consequently, our Louvain implementation achieves a speedup of 6.7×6.7$\times$6.7 × compared to cuVite, and a speedup of 7.3×7.3$\times$7.3 × relative to Rundemanen.